# Learning Physics From Video: Unsupervised Physical Parameter Estimation for Continuous Dynamical Systems

### v1

@@ 4.6 Real-world video evaluation
5 Discussion and Limitations
We present a novel method to perform physical parameter estimation of governing equations. While previous methods in literature do not study phenomena other than motion, we go beyond motion and include a variety of dynamical systems.
We avoid frame prediction, which is a challenging problem but can be used for further visualization. However, this requires a decoder that needs to be designed based on the particular system. A general transformation-agnostic decoder might fail in proper reconstructions. Since our approach does not require decoders, it is applicable to different dynamical systems.
In addition, we examine latent space predictions directly. In contrast, state-of-the-art models do not discuss the accuracy of predicted dynamics compared to the real dynamics [15, 18, 20], but mainly report the frame reconstruction accuracy. However, when the main objective is parameter estimation, reconstruction is simply a tool to define the unsupervised loss, and therefore, the latent space should be analysed closely.
Our model does not distinguish the absolute scale of the variable state, which means the model does not have the means to know that the pendulum goes from $[-90^{\circ},90^{\circ}]$. Yet, thanks to the loss function, the model solves its own metric, ensuring assumptions made in section 3. Baselines implicitly or explicitly do this normalization using the spatial transform, forcing the prediction to be in the pixel metric; instead, all our dynamics depend on the prior distribution assumed by KL-divergence.
Baseline methods are sensitive to the initial value of the parameters $\gamma$. In our method, we also observed that small learning rates with high initial parameters can lead to a local minimum. Therefore, we enforce variance in the learning rate depending on the initial values of the parameters (as explained in Sec. 3, which allow for wider search spaces. Finally, our successful parameter estimation without masks on a real-world video (see Sec. 4.6) is an improvement over the baselines.
Limitations
We used continuous, autonomous differential equations. However, some systems, such as fluids, are described with more complex differential equations. In addition, we need to guarantee that Eq. 8 can be differentiated. Because of the previous reasons, the research in this solution has yet to be explored with an extension of the experiments and more complex use cases, which might include combining dynamics or multiple entities in a scene with independent dynamics.
Due to discretization, performance may vary; for example, small changes in dynamics in the scale dataset were challenging to study because of the resolution of the images.
We avoid mask prediction for the objects in video (see Sec. 4.1). In realistic settings with complex backgrounds, object masks are not necessary as input to train, as real-world experiments showed in section 4.6, while baselines require such masks [18, 15]. However, when the system involves multiple objects interacting, we need to use masks as well, as shown in experiment 4.5. While learning the masks is an interesting solution proposed in [16, 18], their method highly depends on the spatial transformer constrained to roto-translation problems.
Appendix A Appendix / supplemental material
Figure 9: Dataset baseline. It shows the evolution of the spring dynamical system of two MNIST digits over a static CIFAR10 background. Figure edited from [18]
A.1 Dataset Dimentionality
Each uses an image size of $(50\times 50)$ pixels, and the simulated dynamics are the same for the three datasets (Eq. 2) normalized with respect to the image size and maximum pixel intensity. Each dataset consists of 500 training samples with 20 frames, with a final numerical dimensionality of $(samples\times frames\times\#channels\times width\times height)=(500\times 20\times 1\times 50\times 50)$.
A.2 Training Hyperparameters
The experiments and baselines were executed on a GPU NVIDIA 3080. For our model, implemented in PyTorch, the encoder was trained using the Adam optimizer [22]with a learning rate of $1\times e^{-2}$ and the default weight initialization for MLP layers.
A.3 Simulation Details
Here we discuss the details of the dynamics simulation of the experiments in section 4.3. The equation 2 represents and harmonic oscillator with close solution:
$z(t)=Ae^{-\zeta t}cos(\omega t+\phi)$ (14)
$\gamma_{0}=\omega^{2}+\zeta^{2}=4.0016$ (15)
$\gamma_{1}=2\zeta=0.08$ (16)
A.4 Baseline Dataset
In this section, we describe the dataset used to compare the baselines. Both baselines were tested in the dataset published by [18] and also used in [15], the equations of motion used for each system are presented in Eq. 17
$\vec{F}_{ij}=-k(\vec{p}_{i}-\vec{p}_{j})-l\frac{\vec{p}_{i}-\vec{p}_{j}}{|\vec{p}_{i}-\vec{p}_{j}|}$ (17)

### v2

@@ A.7 Delfys75 real-video training details
A.8 Ablation study $dt$
For this experiment, we examine how changing the Euler step $dt$ affects the training algorithm. Note that, due to the properties of Euler’s method, high $dt$ values yield poor approximations, while very small $dt$ values weaken the gradient on the learnable parameter. The result are shown on table 6.
Dataset (Estimated parameter) $dt$ Value
Intensity 0.2 0.08
(Damping factor) 0.4 0.078
$\gamma$ 0.8 0.077
Torricelli 0.1 0.0089
(Flow rate) 0.2 0.0089
$k[\frac{\sqrt{m}}{s}]$ 0.4 0.011
Pendulum 0.1 0.45
(String length) 0.2 0.51
$L[m]$ 0.4 0.48
Table 6: Estimated parameter for various $dt$. Results indicate that it has a small impact on predictions.
A.9 Real video latent space visualization
This section presents two different systems: 1. An LED light video recording with the constant brightness change over time 11, similar to the intensity problems previously studied. 2. A pendulum 10 recording which validates the model in a realistic version of the synthetic dataset. For training these models, no masks were needed compared to baselines [20, 17].
For an LED recording with constant change of brightness over time, 11 shows that the model performs accurately (model predictions in blue) when compared with the ground truth intensity values manually extracted for each frame (red). In Figure 10, we present a more realistic use case where we do not have access to the ground truth or precise manual annotations for each frame. Fortunately, the length of the string $L$ is known to be 120 cm (string length is not visible in the video). The task in this experiment is to learn the value of $L$. Quantitatively, in Figure 10a, the model can reliably estimate the value of the length parameter $L$. Besides, in the latent space Figure 10b, we can see that the model can accurately predict the natural damped oscillations.
Figure 11: Estimating dynamics from a real-world linear intensity change video. At the top are the estimated dynamics from the model (Latent Space), which are compared with a manually extracted ground truth; the plot shows three dots where example frames of the video inputs are shown in the bottom. The plot shows the latent space can accurately estimate the intensity, capturing the global behaviour over time and following the expected dynamics.